# Trapped-ion quantum simulations for condensed-phase chemical dynamics: seeking a quantum advantage

### v3

@@ Introduction
Providing a concrete model on which a quantum advantage can be achieved requires a rigorous study of the cost of classical-digital computation, which is beyond the scope of this Review. Instead, we guide the readers to the state-of-art numerically exact classical-digital simulation results in Fig. 4(a) of Ref. 167 and Sec. V B of Ref. 168. While the models simulated here are not in the form of LVCM in (1), as the bath modes are dissipative (see the Outlook section below for trapped-ion simulations of dissipative bath modes), we expect these results to give a rough sense of the computational cost of numerically exact simulations of LVCMs and similar models that are more complicated than the simple model in (5). Ref. 167 simulated a homogeneous polymer of 20 electronic states with nearest-neighbor electronic coupling $\Delta$. Each electronic state is independently coupled to 2 bath modes (total 40 modes) with reorganization energy 1.25 times larger than the coupling strength between neighboring electronic states ($\lambda/\Delta=1.25$). Simulations that achieve numerical convergence required 60 hours of computation on 2 processors with 8 cores each, although allowing larger errors led to significantly reduced run time. Ref. 168 simulated a similar model with 10 electronic states each coupled to 4 bath modes (total 40 modes) and $\lambda/\Delta=1.5$, which required several days of computation on a 16-core node of a computing cluster. Based on the observation in Fig. 1b that the run time of a numerically exact classical-digital simulation method increases nearly exponentially with $\lambda/\Delta$, it seems possible that a similar model with a few tens of electronic states and bath modes and larger $\lambda/\Delta$ (e.g., $\lambda/\Delta\approx 10$) is challenging for numerically exact simulations, especially when a more complicated connectivity of couplings between the electronic states and bath modes is considered.
@@
In some approximations [53, 54, 231, 232] or analytical regimes [233, 168], a bath consisting of a finite number of damped modes may have the same effects on electronic energy transfer with a bath of infinitely many undamped modes with a continuous (e.g., sum of Lorentzian) spectral density. A bath of continuously distributed modes is different from that described in the LVCM, but is also used widely to study molecule-bath interactions [53, 54, 234, 235, 236, 237, 238, 239]. For trapped ions, directly simulating a finite number of damped modes is more feasible [12, 232, 240] than simulating infinitely many undamped modes, as discussed in the next subsection. Learning appropriate finite-bath approximations for chemical dynamics [232, 233, 168] (or developing approaches to include effectively infinite baths) will be required for the broad applicability of trapped-ion simulations for condensed-phase chemical dynamics.
@@
Two examples illustrate the relative effectiveness of quantum-classical and semiclassical approximations for simulating the LVCM. First, the quantum-classical Liouville equations [101], which approximate open-system quantum dynamics, are exact for the linearly-coupled vibronic dynamics of (1) [250]. Second, in semiclassical theory, the Feynman-Vernon influence functional plays a vital role [251, 252, 253, 254]. Remarkably, the form of this functional is analytic in the limit of a linearly-coupled harmonic bath in (1) [254]. This analytic form may then be used instead of numerically calculating the functional. While the quantum-classical Liouville equations and semiclassical approximations based on the Feynman-Vernon influence functional [70] remain difficult to solve, they are computationally less expensive than direct propagation of the time-dependent Schrodinger equation for simulating the LVCM. This suggests that the LVCMs may not be the best target for attempts to achieve a quantum advantage using analog quantum simulations, as the LVCMs constitute a regime of exceptional classical-digital performance.
@@
Phonon gain also can be similarly simulated by intentionally heating the motional modes, which also requires two kinds of ions [262]. Alternatively, phonon gain can be simulated by averaging over many instances of random stochastic operations interleaved with the usual operations for simulating the LVCM. Here, the random stochastic operations are described by (3) with varying amplitude ($\tilde{\Omega}$), phase ($\varphi$), and/or detuning ($\mu$) of the laser, which is applied to an auxiliary ion that shares the motional mode. The rate of phonon gain determines the range over which the laser parameters are randomly drawn. This method relies on the fact that evolution with respect to the Lindblad master equation can be realized by an average of many evolutions with respect to a closed system consisting of stochastic Hamiltonian terms [263, 264]. Trapping two kinds of ions is not required with this alternative method. Instead, averaging over a large number of analog-simulation runs may be required in order to achieve a sufficiently low uncertainty in simulating the electronic-state population evolution [139].
Second, the second-order coupling between electronic states and bath modes can be simulated [12]. Intuitively, as a spin-dependent force (resonant with the first-order sideband $\omega_{0}\pm\omega_{k}$) can be used in simulating a linear coupling that is proportional to $\hat{\psi}^{\dagger}_{i}\hat{\psi}_{j}(\hat{a}_{k}+\hat{a}_{k}^{\dagger})$, a force resonant with the second-order sideband $\omega_{0}\pm\omega_{k}\pm[\mp]\omega_{l}$ can be used to simulate a second-order coupling that is proportional to $\hat{\psi}^{\dagger}_{i}\hat{\psi}_{j}(\hat{a}_{k}\hat{a}_{l}+\hat{a}_{k}^{\dagger}\hat{a}_{l}^{\dagger})$ [$\hat{\psi}^{\dagger}_{i}\hat{\psi}_{j}(\hat{a}_{k}\hat{a}^{\dagger}_{l}+\hat{a}_{k}^{\dagger}\hat{a}_{l})$]. This was demonstrated in several experiments [265, 213, 266, 267, 268, 269]. A protocol for simultaneously simulating multiple second-order coupling terms is shown in Ref. 270. The challenge in experimental implementations is that the second-order sideband interactions are typically an order of magnitude weaker than the first-order interactions. Thus, longer coherence times of the experimental system may be required to simulate the dynamics with second-order vibronic couplings accurately. Similar arguments apply to simulating higher-order vibronic couplings.

### v5

@@ Introduction
Providing a concrete model on which a quantum advantage can be achieved requires a rigorous study of the cost of classical-digital computation, which is beyond the scope of this Review. Instead, we guide the readers to the state-of-art numerically exact classical-digital simulation results in Fig. 4(a) of Ref. 75 and Sec. V B of Ref. 168. While the models simulated here are not in the form of LVCM in (1), as the bath modes are dissipative (see the Outlook section below for trapped-ion simulations of dissipative bath modes), we expect these results to give a rough sense of the computational cost of numerically exact simulations of LVCMs and similar models that are more complicated than the simple model in (5). Ref. 75 simulated a homogeneous polymer of 20 electronic states with nearest-neighbor electronic coupling $\Delta$. Each electronic state is independently coupled to 2 bath modes (total 40 modes) with reorganization energy 1.25 times larger than the coupling strength between neighboring electronic states ($\lambda/\Delta=1.25$). Simulations that achieve numerical convergence required 60 hours of computation on 2 processors with 8 cores each, although allowing larger errors led to significantly reduced run time. Ref. 168 simulated a similar model with 10 electronic states each coupled to 4 bath modes (total 40 modes) and $\lambda/\Delta=1.5$, which required several days of computation on a 16-core node of a computing cluster. Based on the observation in Fig. 1b that the run time of a numerically exact classical-digital simulation method increases nearly exponentially with $\lambda/\Delta$, it seems possible that a similar model with a few tens of electronic states and bath modes and larger $\lambda/\Delta$ (e.g., $\lambda/\Delta\approx 10$) is challenging for numerically exact simulations, especially when a more complicated connectivity of couplings between the electronic states and bath modes is considered.
@@
In some approximations [54, 55, 231, 232] or analytical regimes [233, 168], a bath consisting of a finite number of damped modes may have the same effects on electronic energy transfer with a bath of infinitely many undamped modes with a continuous (e.g., sum of Lorentzian) spectral density. A bath of continuously distributed modes is different from that described in the LVCM, but is also used widely to study molecule-bath interactions [54, 55, 234, 92, 235, 236, 237]. For trapped ions, directly simulating a finite number of damped modes is more feasible [12, 232, 238] than simulating infinitely many undamped modes, as discussed in the next subsection. Learning appropriate finite-bath approximations for chemical dynamics [232, 233, 168] (or developing approaches to include effectively infinite baths) will be required for the broad applicability of trapped-ion simulations for condensed-phase chemical dynamics.
@@
Two results illustrate the relative effectiveness of quantum-classical and semiclassical approximations for simulating the LVCM. First, the quantum-classical Liouville equations [104] are used to approximate open-system quantum dynamics. This approach becomes exact for the linearly-coupled vibronic dynamics of (1) [248]. Second, more generally in semiclassical theory, the Feynman-Vernon influence functional plays a vital role [249, 250, 251, 252]. This functional encodes the influence of a bath on the dynamics of the system of interest as a function of the trajectory of the system of interest only [250], including any influence that encodes memory of previous states of the system of interest. Remarkably, the form of this functional is analytic in the limit of a linearly-coupled harmonic bath in (1) [253, 252]. This analytic form may then be used instead of numerically calculating the functional. While the quantum-classical Liouville equations and semiclassical approximations based on the Feynman-Vernon influence functional [71] remain difficult to solve, they are computationally less expensive than direct propagation of the time-dependent Schrodinger equation for simulating the LVCM. This suggests that the LVCMs may not be the best target for attempts to achieve a quantum advantage using analog quantum simulations, as the LVCMs constitute a regime of exceptional classical-digital performance.
@@
Phonon gain also can be similarly simulated by intentionally heating the motional modes, which also requires two kinds of ions [261]. Alternatively, phonon gain can be simulated by averaging over many instances of random stochastic operations interleaved with the usual operations for simulating the LVCM. Here, the random stochastic operations are described by (3) with varying amplitude ($\tilde{\Omega}$), phase ($\varphi$), and/or detuning ($\mu$) of the laser, which is applied to an auxiliary ion that shares the motional mode. The rate of phonon gain determines the range over which the laser parameters are randomly drawn. This method relies on the fact that evolution of an open quantum system with respect to the Lindblad master equation can be realized by an average of many evolutions of closed systems, each consisting of stochastic Hamiltonian terms [262, 263]. Trapping two kinds of ions is not required with this alternative method. Instead, averaging over a large number of analog-simulation runs may be required in order to achieve a sufficiently low uncertainty in simulating the electronic-state population evolution [142].
Second, the second-order coupling between electronic states and bath modes can be simulated [12]. Intuitively, as a spin-dependent force (resonant with the first-order sideband $\omega_{0}\pm\omega_{k}$) can be used in simulating a linear coupling that is proportional to $\hat{\psi}^{\dagger}_{i}\hat{\psi}_{j}(\hat{a}_{k}+\hat{a}_{k}^{\dagger})$, a force resonant with the second-order sideband $\omega_{0}\pm\omega_{k}\pm[\mp]\omega_{l}$ can be used to simulate a second-order coupling that is proportional to $\hat{\psi}^{\dagger}_{i}\hat{\psi}_{j}(\hat{a}_{k}\hat{a}_{l}+\hat{a}_{k}^{\dagger}\hat{a}_{l}^{\dagger})$ [$\hat{\psi}^{\dagger}_{i}\hat{\psi}_{j}(\hat{a}_{k}\hat{a}^{\dagger}_{l}+\hat{a}_{k}^{\dagger}\hat{a}_{l})$]. This was demonstrated in several experiments [264, 213, 265, 266, 267, 268]. A protocol for simultaneously simulating multiple second-order coupling terms is shown in Ref. 269. The challenge in experimental implementations is that the second-order sideband interactions are typically an order of magnitude weaker than the first-order interactions. Thus, longer coherence times of the experimental system may be required to simulate the dynamics with second-order vibronic couplings accurately. Similar arguments apply to simulating higher-order vibronic couplings.